OpenStudy (anonymous):

Help! There are 5 black males and 6 black females. There are 8 white males and 4 white females. Suppose you choose 3 randomly. What is the probability that you have atleast a black and atleast a female within the 3?

OpenStudy (anonymous):

@mathmate @jim_thompson5910 @completeidiot

OpenStudy (anonymous):

n(S)=C(23,3)=1771 \[E=B \cap F \\ E'=B' \cup F'\] \[n(B)=C(11,3)=165 \\ n(F)=C(10,3)=120\] \[n(B \cap F)=C(6,3)=20 \\ n(E')=C(11,3)+C(10,3)-C(6,3)=265 \\ n(E)=n(S)-n(E')=1771-265=1506\] That's what I have, but it doesn't seem right!

OpenStudy (mathmate):

Have you done Binomial distribution?

OpenStudy (anonymous):

I don't believe so unless they have a different name, but if they would help solve this problem, then I'm up for learning about them.

OpenStudy (mathmate):

I would separate the problem into two, one for M/F, and the other one B/W. Calculate P(F) for 3 picks, and P(B) for 3 picks. Then we can find \( P(F) \cup P(B) \).

OpenStudy (anonymous):

Wouldn't it be the intersection because you need one of each and not just one which you would have in a union?

OpenStudy (mathmate):

You're right, my bad!

OpenStudy (mathmate):

In a binomial experiment with 3 trials, N=3 (trials) p=10/23 (female) q=1-p=13/23 (male) P(all male)=C(3,3)p^0q^3=1*(10/23)^0*(13/23)^3=.1806 P(at least one female)=1-.1806=.8194 Similarly for B/W.

OpenStudy (anonymous):

Thanks!

OpenStudy (mathmate):

You're welcome! :)